VIBRATIONAL ANHARMONICITY IN ETHYLENE, METHYL FLUORIDE AND DICHLOROMETHANE: AN EXPLORATION USING GAUSSIAN 03

\underline{NORMAN C. CRAIG}, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; MARK M. LAW, Department of Chemistry, University of Aberdeen, Meston Walk, Aberdeen, AB24 3UE, U.K.; DONALD C. MCKEAN, School of Chemistry, University of Edinburgh, West Mains Road, Edinburgh, EH9 3JJ, U.K..

This survey was undertaken (a) to test the suitability of the program Gaussian 03 (G03) for anharmonicity calculations: (b) to compare results for different models (B3LYP, MP2) and basis sets (6-311++G**, cc-pVTZ): (c) to see if anharmonicity corrections to harmonic frequencies were identical for different types of C-H stretching vibration. Results to be presented show that the version of G03 available was inadequate to detect familiar instances of Fermi resonance. For C_{3v}-type molecules G03 yielded anharmonicity constants x_{i,j} unusable by the experimentalist. In CH_{2}Cl_{2} the antisymmetric C-H stretching frequency is associated with a larger anharmonicity correction than is the case for the C-H stretch, in a situation where Fermi resonance is negligible. Calculated values of x_{1,1}, x_{1,6} and x_{6,6} for CH_{2}Cl_{2} agree excellently with experimental data. In ethylene a marked effect of basis set was found in MP2 calculations for the out-of-plane bending modes, linked to the presence or absence of f functions.